Van Frank Telecommunications has a patent on a cellular transmission process. The company has amortized the patent on a straight-line basis since 2014, when it was acquired at a cost of $11.7 million at the beginning of that year. Due to rapid technological advances in the industry, management decided that the patent would benefit the company over a total of six years rather than the nine-year life being used to amortize its cost. The decision was made at the beginning of 2018.

Required:
Prepare the year-end journal entry for patent amortization in 2018. No amortization was recorded during the year. (If no entry is required for a transaction/event, select "No journal entry required" in the first account field. Enter your answer in millions rounded to 1 decimal place (i.e., 5,500,000 should be entered as 5.5)).
Answer #1

Previous amortization per year = 11.7/9 =1.3 per year

Book value of patent at the beginning of 2018

= 11.7 - (1.3*4 years) = 6.5

New amortization per year = 6.5/(6-4) = 3.25

Amortization expense 3.25
Patents 3.25